1. Solve the given differential equation by undetermined coefficients.
y'' + 3y = −48x^2 (e^3x).

y(x) =____


2. Find a particular solution of the given differential equation. Use a CAS as an aid in carrying out differentiations, simplifications, and algebra.

y^(4) + 2y'' + y = 1 cos x − 3x sin x

yp = _____
1
Expert's answer
2020-06-30T16:31:04-0400


1.y'' + 3y = −48x2e3x.

Solution:

y'' + 3y=0

"\lambda"2+3 =0

"\lambda" 1=-i31/2, "\lambda" 2=i31/2.

y0(x)=c1cos(31/2 x)+c2sin(31/2x),

yp(x)=(Ax2+Bx+C)e3x

yp'(x)=(3Ax2+(3B+2A)x+3C+B)e3x

yp''(x)=(9Ax2+(9B+12A)x+2A+6B+9C)e3x

x2e3x:3A+9A=-48,

xe3x:3B+9B+12A=0,

e3x:3C+2A+6B+9C=0.

=>A=-4,B=4,C=-4/3.

yp(x)=(-4x2+4x-4/3)e3x,

y(x)=y0(x)+yp(x)

Answer:

y(x)=c1cos(31/2 x)+c2sin(31/2x)+(-4x2+4x-4/3)e3x

2.y(4) + 2y'' + y = cos x − 3x sin x 

Solution:

y(4) + 2y'' + y =0

"\lambda" 4+2"\lambda" 2+1=0

("\lambda" 2+1)2=0

"\lambda" 1=-i, k=2,

"\lambda"2 =i, k=2

yp(x)=(Ax3+Bx2)cos(x)+(Cx3+Dx2)sin(x)

yp'(x)=...,yp''(x)=...yp'''(x)=...,y(4)(x)=...

yp(4) + 2yp'' + yp=cos(x)(-24Ax-8B+24C)+sin(x)(-24A-24Cx-8D)

-24A=0

-8B+24C=1

-24C=3

-24A-8D=0

=>A=0,B=1/4, C=1/8, D=0

Answer:

yp(x)=x2/4cos(x)+x3/8sin(x)